Prev Next

SW&IT Meeting Minutes 2020-03-16

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

News from management

  • COVID-19 and remote
    • CV email (Re: COVID-19 and LBTO), please read in its entirety and carefully ... on the more practical side, extracted quote:

Whatever you do:

- You must let your supervisor know about you whereabouts!

- If you need to bring home equipment beyond your laptop to make your work efficient, such as monitors or desktops, you can but you must do the following:

Take pictures of each piece of equipment and its LBT tag (or model/serial number factory tag if it is not LBT tagged) and just email them to admin@lbto.org with an appropriate subject line...

Remember that all face-to-face meetings with more than 4 people must be virtual (Zoom)....

Your supervisors will address issues specific to your work as deemed necessary.

  • zoom3 allocated for sciops open channel

Previous Action Items

OwnerSorted ascending Action Due Date Completion Date
DE timecapsule obsdev 2020-02-12  
DE oVirt for PEPSI 2020-02-12  
DE NAS synology status    
IS stolen laptop    

Open topics

  • COVID-19/remote challenges - share forseen acivities for next 2 weeks
  • Tucson CR cooling
  • SSD fiber new bundle inside new PEPSI chain - regular follow up
  • new OWLs vs zoom-polycom (to be discussed at ROWG#4)

Team reports

YZ

  • Merged recent LBC and MODS scripts to github.
  • Organizing AO Strehl codes (to be merged to github as well).
  • Working on some improvements to the MODS data cleaning script.
  • Working with Olga on 64-bit MODS programs (make them work on robs).
  • Moved back to 5th floor.
  • Set up an LBTO phone.
  • Working remote mostly for the time being.

DE

  • Polycom-->Zoom bridge setup with dedicated meetiing ID
  • KS swgdev01 F30, install Matlab
  • Mac mini further delay, slated to ship 19 March
    • Stephen located a 4TB Time Capsule
    • Have been serially wiping drive to test for bad blocks; looks good so far
  • Synology is in pretty good shape. I have been working on getting prefail drives replaced.
  • VMHost arrays back to full parity
  • Added CPU to virtual abell; seems to have stopped the regular CPU limit warnings

MB

  • Last week
    • OC, WPC, Management
    • DMS from matlab (J.Crass for iLocater)
    • ALTA follow-up, predictions when iced
    • Covid-19 and ROWG related activities
  • This week
    • Covid-19, remote new arrangement, and making sure it works
    • ROWG related activitie

XZ

  • Last week
    • DX AdSec troubleshoting and tests on the mountain
      • cool the clean room
      • test the unit in the mid-night and early morning
      • dicussion with Guido, Mario and Roberto
      • read dsp memory
      • fastlink test modification
    • Support DX AdSec from Tucson
      • thermal test
      • read dsp memory
      • dicussion with Guido, Mario
    • SX AdSec issue
  • This week
    • SX AdSec issue
    • write python routine to read the telemetry file and plot
    • Power Backplane?

IS

Last week:
  • PIT
    • Starting working on migrating changes from the general PIT to the Italian PIT. A little bit of a pain because both start differing in structure. I.e. the Italian PIT doesn't have a binocular concept so it is not as easy as migrating commits.
    • working on the login part of the PIT admin tool backend.
  • LUCI
    • Talked to Dave Thompson about how to make LUCI engineering easier for engineers and us. Ideally come up with scripts that work with what we currently have, (by might not be possible) and create tools that minimize interaction with software.
  • OT
    • Created a new OT. The operations teams have been swamp and couldn't finish the OT testing on time. So now I'm creating a small bug release on the previous version the 2019B.1.1.0 to 2019B.1.2.0

This week:
  • OT
    • Hopefully release latest OT. Have to setup QA to current version
  • PIT
    • Italian PIT
    • PIT admin tool
  • LUCI
    • Work on sketching engineering/test/support scripts

SH

  • move back to fifth floor
  • a little bit of DMS
  • remote observing work
  • VPN for people
  • DNS changes
  • rm507 stability issues

next:

  • ovirt?
  • ROWG
  • LUCI fonts

CH

Last week:
  • Vacation

This week:
  • LBTI client, add alarms to AHS.
Topic revision: r10 - 16 Mar 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback